# Precision Determination of Scintillation Screen Parameters from Annual Modulation Measurement of Pulsar Scintillation Arc Curvature with the FAST Telescope

### v1

@@ Abstract
software: AstroPy (Astropy Collaboration et al., 2013, 2018, 2022), corner (Foreman-Mackey, 2016), DSPSR (van Straten & Bailes, 2011), emcee (Foreman-Mackey et al., 2013), lmfit (Newville et al., 2016), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007) and NumPy (Harris et al., 2020)
1 Introduction
@@
Pulsars are highly magnetized, rotating neutron stars that emit beams of radio waves from their magnetic poles (Hewish et al., 1968; Gold, 1968). Inhomogeneities in the electron density of the ionized interstellar medium (IISM) scatter a pulsar’s radio waves as they propagate, leading to multi-path interference. As the Earth, the pulsar, and the IISM move relative to each other, the interference conditions change with time. Consequently, the pulsar’s radio signal intensity is modulated in time and frequency, a phenomenon known as interstellar scintillation (ISS) (Scheuer, 1968). Observations and analysis of pulsar scintillation provide a powerful probe of the IISM. The scattering occurs on spatial scales of only about an astronomical unit (AU) (Rickett, 1990), making ISS an effective tool for investigating extremely small-scale structures in the IISM.
@@
Subsequent observations have shown that interstellar scintillation is concentrated in discrete regions along the line of sight referred to as “scintillation screens”, and that the scintillation screens are prevalent in the IISM. A survey conducted using the Green Bank Telescope (GBT) and the Arecibo Observatory (Stinebring et al., 2022) detected definite or probable scintillation screens in 19 out of 22 pulsars with low to moderate dispersion measure. Wu et al. (2022) conducted the first sizeable census of diffractive pulsar scintillation at low frequencies (120 - 180 MHz) using the Low-Frequency Array (LOFAR), detecting scintillation screens in 9 out of the 15 pulsars that allow for a full determination of the scintillation properties. Main et al. (2023) observed 107 pulsars with interstellar scintillation screens using the MeerKAT Thousand Pulsar Array Program, concluding that these screens appear to be ubiquitous in clean, high signal-to-noise observations. High-sensitivity measurements have revealed multiple distinct arcs along the line of sight to individual pulsars. For instance, Ocker et al. (2024) observed at least 9 separate screens toward B1929+10 with the Five-hundred-meter Aperture Spherical radio Telescope (FAST), Reardon et al. (2025) used the MeerKAT radio telescope to discover 25 screens in the direction of the millisecond pulsar PSR J0437-4715.
@@
Scintillation screens can be associated with various interstellar structures, including H II regions (Mall et al., 2022), bow shocks (Reardon et al., 2025), and supernova remnants (Yao et al., 2021). Establishing connections between scintillation screens and these structures can reveal the origin of the scintillation screens while simultaneously providing insights into both the structures themselves and the physical properties of the associated pulsars. For instance, Yao et al. (2021) discovered a scintillation screen associated with the supernova remnant, and by utilizing this screen as a probe, they obtained the first evidence that the three-dimensional direction of motion in a young pulsar aligns with its rotation axis. Similarly, Reardon et al. (2025) linked four scintillation screens to the bow shock, and through analysis of these screens, they revealed details about the bow shock structure as well as the pulsar’s radial velocity.
@@
However, in most cases, the astrophysical associations of scintillation screens are still unknown. Also under debate is the physical nature of the microscopic electron density fluctuations that generate the scintillation. There exists a promising proposal that scintillation screens are related to reconnection sheets - dissipative regions of magnetohydrodynamic (MHD) turbulence in the IISM (Goldreich & Sridhar, 2006; Pen & King, 2012; Pen & Levin, 2014; Liu et al., 2016; Simard & Pen, 2018). A major impediment to understanding the nature of scintillation screens and discovering their potential macroscopic associations is the uncertainty in their distances. So far, most observed scintillation screens still do not have well-determined distances.
@@
In early times, the curvature of a single scintillation arc measurement is used to determine the distance to the screen (e.g., Putney & Stinebring 2006). Such distance determination neglects the possible transverse velocity of the scintillation screen and assumes that the scattering is isotropic in the scintillation screen. Later observations have found a high degree of anisotropy in the scattering in some scintillation screens (e.g., Brisken et al. 2010, Sprenger et al. 2022, Stinebring et al. 2022). Considering this high degree of anisotropy, the parabolic arc curvature depends on three parameters of the screen: the distance from the observer, the main scattering direction, and the velocity projected along the main scattering direction. VLBI observations of pulsar scintillation are effective in lifting the parameter degeneracy by providing phase information (Brisken et al., 2010), but are very costly to perform. Apart from VLBI, the only feasible method to measure these screen parameters is to use the annual modulation of parabolic arc curvature. By observing the scintillation pattern of a pulsar at multiple epochs throughout an observing year, one can verify the annual modulation nature of the parabolic arc curvature variation, and precisely determine these screen parameters (e.g., Reardon et al. 2020; McKee et al. 2022; Mall et al. 2022; Xu et al. 2023; Liu et al. 2023).
@@
This paper is organized as follows: Section 2 discusses the theory of the annual modulation of scintillation arc curvature; Section 3 describes the observational data and data processing methods; Section 4 presents the results; Section 5 provides a summary of our findings and an outlook for future research.
2 Theory of annual modulation
@@
The scintillation arc curvature depends on the observing frequency $\nu$, the pulsar distance $\text{d}_{\text{psr}}$, the fractional distance s of the scintillation screen to the pulsar, and the projected relative velocity of the screen with respect to the pulsar-observer line of sight $\text{V}_{\text{sr}}$ (see e.g. Brisken et al. 2010),
$\eta_{\nu}=\frac{c\text{d}_{\text{psr}}\text{s}(1-\text{s})}{2\nu^{2}\text{V}_{\text{sr}}^{2}}\,,$ (1)
@@
$\text{V}_{\text{sr}}=\left|\text{s}\mathbf{V_{\text{Earth}}}+(1-\text{s})\mathbf{V_{\text{psr}}}-\text{V}_{\text{screen}}\mathbf{n_{\text{screen}}}\right|.$ (2)
Here, all velocities are 2D transverse velocities with respect to the pulsar line of sight. $\mathbf{V_{\text{Earth}}}$ and $\mathbf{V_{\text{psr}}}$ are the Earth’s and the pulsar’s velocities, respectively. $\text{V}_{\text{screen}}$ is the velocity of the screen and $\mathbf{n_{\text{screen}}}$ is its 2D orientation vector in the sky.
@@
$\text{V}_{\text{sr}}=\left[\text{s}\mathbf{V_{\text{Earth}}}+(1-\text{s})\mathbf{V_{\text{psr}}}\right]\cdot\mathbf{n_{\text{screen}}}-\text{V}_{\text{screen}}.$ (3)
In this case, $\mathbf{n_{\text{screen}}}$ stands for the main scattering direction and $\text{V}_{\text{screen}}$ is the screen velocity along this direction. We shall limit $\mathbf{n_{\text{screen}}}$ to be within [0, 180) degrees of a specified direction.
@@
$f_{\eta}\equiv\frac{\sqrt{2}\text{V}_{\text{sr}}}{\sqrt{s(1-s)\text{d}_{\text{psr}}}}.$ (5)
@@
Given the distance and the proper motion of the pulsar, we are left with three unknown parameters s, $\text{V}_{\text{screen}}$, and $\mathbf{n_{\text{screen}}}$. They are to be fitted from the annual modulation of the measurements $|f_{\eta}|(t)$.
2.1 Analytical understanding of the parameter dependence
2.1.1 1D scattering
Let us first consider the simpler case of 1D scattering, i.e., extremely anisotropic scattering. Approximating the Earth’s orbit as circular, $f_{\eta}(t)$ has a sinusoidal shape that is characterized by three quantities: the modulation amplitude $\text{A}_{\text{f}}\equiv(f_{\eta\text{, max}}-f_{\eta\text{, min}})/2$, the mean offset $\bar{f_{\eta}}$, and the yearly time $t_{\text{max}}$ when $f_{\eta}(t)$ reaches its maximum. They are dependent on the three parameters as follows.
When projected onto the direction of the pulsar, the Earth’s velocity vector $\mathbf{V_{\text{Earth}}}(t)$ forms an ellipse around the year (see Figure 1). This velocity ellipse is further projected onto the direction of $\mathbf{n_{\text{screen}}}$ to obtain $\text{V}_{\text{proj}}\equiv\mathbf{V_{\text{Earth}}}\cdot\mathbf{n_{\text{screen}}}$. The modulation amplitude $\text{A}_{\text{f}}$ is set by the maximum of $\text{V}_{\text{proj}}$,
$\text{A}_{\text{f}}=\sqrt{\frac{2\text{s}}{(1-\text{s})}}\frac{\text{V}_{\text{proj,max}}}{\sqrt{\text{d}_{\text{psr}}}},$ (6)
and the maximum of $f_{\eta}(t)$ is reached at time $t_{\text{max}}$ when $|\text{V}_{\text{proj}}|=\text{V}_{\text{proj,max}}$. Therefore, from $t_{\text{max}}$ one can determine the orientation $\mathbf{n_{\text{screen}}}$. Given $\mathbf{n_{\text{screen}}}$, one can compute $\text{V}_{\text{proj,max}}$ and obtain s from the measurement of the modulation amplitude $\text{A}_{\text{f}}$.
@@ 2.1 Analytical understanding of the parameter dependence
$\bar{f_{\eta}}=\left[1-\frac{1}{1-\text{s}}\frac{\text{V}_{\text{screen}}}{\mathbf{V_{\text{psr}}}\cdot\mathbf{n_{\text{screen}}}}\right]\sqrt{\frac{2(1-\text{s})}{\text{s}}}\frac{\mathbf{V_{\text{psr}}}\cdot\mathbf{n_{\text{screen}}}}{{\sqrt{\text{d}_{\text{psr}}}}}.$ (7)
It depends on all three parameters. When the other two parameters are determined from $t_{\text{max}}$ and $\text{A}_{\text{f}}$, one can further determine $\text{V}_{\text{screen}}$ from $\bar{f_{\eta}}$.
Since a physical screen velocity is typically much smaller than the pulsar proper motion velocity, one generally has $\text{V}_{\text{screen}}/(\mathbf{V_{\text{psr}}}\cdot\mathbf{n_{\text{screen}}})/(1-\text{s})\ll 1$, i.e.
$\bar{f_{n}}\approx\sqrt{\frac{2(1-\text{s})}{\text{s}}}\frac{\mathbf{V_{\text{psr}}}\cdot\mathbf{n_{\text{screen}}}}{{\sqrt{\text{d}_{\text{psr}}}}}.$ (8)
@@
$\text{A}_{\text{f}}\bar{f_{\text{n}}}\approx\frac{2\text{V}_{\text{proj,max}}\mathbf{V_{\text{psr}}}\cdot\mathbf{n_{\text{screen}}}}{\text{d}_{\text{psr}}}\,,$ (9)
@@
2.1.2 2D isotropic scattering
The full geometry of a general 2D scattering screen cannot be constrained by the curvature measurements. It requires full modeling of the distribution of scattered power along and within a parabolic arc (Reardon et al., 2020), and is subject to degeneracies. The ideal case of 2D isotropic scattering, however, can be fully captured with a three-parameter model similar to the case of 1D scattering. In the early scintillation studies, scattering was thought to be produced by statistically isotropic ISM turbulence, and thus a 2D isotropic scattering was widely used. Deviation of the shape of $f_{\eta}(t)$ from a sinusoidal shape can be a sign of 2D scattering. For 2D isotropic scattering, the shape of $f_{\eta}(t)$ is in general not symmetric around the mean, but this asymmetric deviation from the sinusoidal shape is only strong when the velocity modulation by Earth is comparable to the yearly average of $\text{V}_{\text{sr}}/\text{s}$.
Apart from this strong-modulation case, the analytical framework for 1D scattering still holds approximately. The only difference is the projection of $\mathbf{V_{\text{Earth}}}$ to $\text{V}_{\text{proj}}$ that should be performed along the direction of $(1-\text{s})\mathbf{V_{\text{psr}}}-\text{V}_{\text{screen}}\mathbf{n_{\text{screen}}}$ instead of $\mathbf{n_{\text{screen}}}$.
3 Observation and Data Processing
3.1 Observation
@@
Table 1: The observed pulsars and their properties. The properties of PSR B1237+25 and PSR B2021+51 were obtained from measurements using the NRAO Very Long Baseline Array (Brisken et al., 2002). The proper motion and the location of PSR B1842+14 were obtained from a program of observations carried out at Jodrell Bank to measure the proper motions of pulsars (Harrison et al., 1993). The DM of PSR B1842+14 was obtained from The Thousand-Pulsar-Array program on MeerKAT (Keith et al., 2024). Yao et al. (2020), using the YMW16 model (Yao et al., 2017), adopted an uncertainty of 40% of the distance (1.7 kpc) as a rough estimate for the one-sigma error of the DM-distance to PSR B1842+14. Therefore, we also adopt a distance of 1.7 ± 0.7 kpc for PSR B1842+14. The $\text{S}_{1400}$ of PSR B1842+14 was also obtained from the research by Yao et al. (2020).
Pulsar RAJ (hms) DECJ (dms) PMRA (mas yr-1) PMDEC (mas yr-1) Parallax (mas) or Distance (kpc) $\text{S}_{1400}$ (mJy) DM ($\text{cm}^{-3}$ pc)
B1237+25 12:39:40.36 +24:53:50.01 $-106.82\pm 0.17$ $49.92\pm 0.18$ $1.16\pm 0.08$ (mas) 20 9.28
B1842+14 18:44:54.88 +14:54:13.89 $-9\pm 10$ $45\pm 6$ $1.7\pm 0.7$ (kpc) 1.8 41.49
B2021+51 20:22:49.87 +51:54:50.39 $-5.23\pm 0.17$ $11.54\pm 0.28$ $0.50\pm 0.07$ (mas) 27 22.58
@@
3.2 Dynamic Spectra
We used DSPSR  [^1] https://psrchive.sourceforge.net/ (van Straten & Bailes, 2011) to fold the data in each frequency channel, with a sub-integration time of approximately 5 s, and then summed the horizontal and vertical polarizations to obtain the total intensity I. Next, we computed the average pulse profile in order to determine the on-pulse and off-pulse regions. We smoothed this average pulse profile. The region below the median of the smoothed profile was identified as the off-pulse region, and the region above 10% of the peak amplitude with respect to the off-level was marked as the on-pulse region. For each sub-integration and frequency channel, we calculated the mean intensity in the on-pulse region minus the mean intensity in the off-pulse region. The result of this calculation is the dynamic spectrum.
We used the level of the off-pulse signal to identify the radio-frequency interference (RFI). After some experimenting, we set 1.5 times the median with respect to the minimum as the threshold for RFI marking.
Next, we applied Savitzky-Golay filters (Savitzky & Golay, 1964) to smooth the dynamic spectra. Finally, we applied Wiener filters (Lin et al., 2021) to the dynamic spectra, filling the masked regions with the expected signal values obtained from the filter.

### v2

@@ Abstract
software: AstroPy (Astropy Collaboration et al., 2013, 2018, 2022), corner (Foreman-Mackey, 2016), DSPSR (van Straten and Bailes, 2011), emcee (Foreman-Mackey et al., 2013), lmfit (Newville et al., 2016), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007) and NumPy (Harris et al., 2020)
I Introduction
@@
Pulsars are highly magnetized, rotating neutron stars that emit beams of radio waves from their magnetic poles (Hewish et al., 1968; Gold, 1968). Inhomogeneities in the electron density of the ionized interstellar medium (IISM) scatter a pulsar’s radio waves as they propagate, leading to multi-path interference. As the Earth, the pulsar, and the IISM move relative to each other, the interference conditions change with time. Consequently, the pulsar’s radio signal intensity is modulated in time and frequency, a phenomenon known as interstellar scintillation (ISS) (Scheuer, 1968). Observations and analysis of pulsar scintillation provide a powerful probe of the IISM. Sub-structure from scintillation arcs and arclets implies spatial scales even smaller than 1 AU by a few orders of magnitude (Rickett, 1990), making ISS an effective tool for investigating extremely small-scale structures in the IISM.
@@
Subsequent observations have shown that interstellar scintillation is concentrated in discrete regions along the line of sight referred to as “scintillation screens”, and that the scintillation screens are prevalent in the IISM. A survey conducted using the Green Bank Telescope (GBT) and the Arecibo Observatory (Stinebring et al., 2022) detected definite or probable scintillation screens in 19 out of 22 pulsars with low to moderate dispersion measure. Wu et al. (2022) conducted the first sizeable census of diffractive pulsar scintillation at low frequencies (120 - 180 MHz) using the Low-Frequency Array (LOFAR), detecting scintillation screens in 9 out of the 15 pulsars that allow for a full determination of the scintillation properties. Main et al. (2022) observed 107 pulsars with interstellar scintillation screens using the MeerKAT Thousand Pulsar Array, concluding that these screens appear to be ubiquitous in clean, high signal-to-noise observations. High-sensitivity measurements have revealed multiple distinct arcs along the line of sight to individual pulsars. For instance, Ocker et al. (2024) observed at least 9 separate screens toward B1929+10 with the Five-hundred-meter Aperture Spherical radio Telescope (FAST), Reardon et al. (2025) used the MeerKAT radio telescope to discover 25 screens in the direction of the millisecond pulsar PSR J0437-4715.
@@
Scintillation screens can be associated with various interstellar structures, including H II regions (Mall et al., 2022), bow shocks (Reardon et al., 2025), and supernova remnants (Yao et al., 2021). Establishing connections between scintillation screens and these structures can reveal the origin of the scintillation screens while simultaneously providing insights into both the structures themselves and the physical properties of the associated pulsars. For instance, Yao et al. (2021) discovered a scintillation screen associated with the supernova remnant, and by utilizing this screen as a probe, they obtained the first evidence that the three-dimensional direction of motion in a young pulsar aligns with its rotation axis. Additionally, scattering associated with the Crab pulsar has been extensively studied since the pulsar’s discovery (Staelin and Reifenstein, 1968), with detailed characterization of nebular scattering (e.g., Cordes et al. 2004, McKee et al. 2018, Main et al. 2021, Lin et al. 2023, Serafin Nadeau et al. 2024). Another example is that Reardon et al. (2025) linked four scintillation screens to the bow shock, and through analysis of these screens, they revealed details about the bow shock structure as well as the pulsar’s radial velocity.
@@
However, in most cases, the astrophysical associations of scintillation screens are still unknown. Also under debate is the physical nature of the microscopic electron density fluctuations that generate the scintillation. There exists a promising proposal that scintillation screens are related to reconnection sheets - dissipative regions of magnetohydrodynamic (MHD) turbulence in the IISM (Goldreich and Sridhar, 2006; Pen and King, 2012; Pen and Levin, 2014; Liu et al., 2016; Simard and Pen, 2018). A major impediment to understanding the nature of scintillation screens and discovering their potential macroscopic associations is the uncertainty in their distances. So far, most observed scintillation screens still do not have well-determined distances.
@@
In early times, the curvature of a single scintillation arc measurement was used to determine the distance to the screen (e.g., Putney and Stinebring 2006). Such distance determination neglects the possible transverse velocity of the scintillation screen and assumes that the scattering is isotropic in the scintillation screen. Later observations have found a high degree of anisotropy in the scattering in some scintillation screens (e.g., Brisken et al. 2010, Sprenger et al. 2022, Stinebring et al. 2022). Considering this high degree of anisotropy, the parabolic arc curvature depends on three parameters of the screen: the distance from the observer, the main scattering direction, and the velocity projected along the main scattering direction. VLBI observations of pulsar scintillation allow for measurement of the scattered brightness distribution with unprecedented precision, revealing fine-scale structures in the scintillation screen, and effectively lifting the parameter degeneracy by providing phase information (Brisken et al. 2010; Stock et al. 2025).
Apart from VLBI, the only feasible method to measure these screen parameters is to use the annual modulation of parabolic arc curvature. By observing the scintillation arc curvature of a pulsar at multiple epochs throughout an observing year, one can determine the main scattering direction by measuring how Earth’s orbital motion projects onto the scattering axis, then determine the screen distance from the variation in arc curvature combined with the pulsar distance, and finally determine the screen velocity from the mean arc curvature combined with the pulsar distance and proper motion (e.g., Reardon et al. 2020; McKee et al. 2022; Mall et al. 2022; Xu et al. 2023; Liu et al. 2023).
@@
This paper is organized as follows: Section II discusses the theory of the annual modulation of scintillation arc curvature; Section III describes the observational data and data processing methods; Section IV presents the derived parameters of the scintillation screens and discusses their distribution and potential sources; Section V provides a summary of our findings and an outlook for future research.
II Theory of annual modulation
@@
The scintillation arc curvature depends on the observing frequency $\nu$, the pulsar distance $d_{\text{psr}}$, the fractional distance $s$ of the scintillation screen to the pulsar, and the projected relative velocity of the screen with respect to the pulsar-observer line of sight $v_{\text{sr}}$ (see e.g. Brisken et al. 2010),
$\eta_{\nu}=\frac{cd_{\text{psr}}s(1-s)}{2\nu^{2}v_{\text{sr}}^{2}}\,$, (1)
@@
$v_{\text{sr}}=\left|s\mathbf{v}_{\text{Earth}}+(1-s)\mathbf{v}_{\text{psr}}-v_{\text{screen}}\mathbf{n}_{\text{screen}}\right|.$ (2)
Here, all velocities are 2D transverse velocities with respect to the pulsar line of sight. $\mathbf{v}_{\text{Earth}}$ and $\mathbf{v}_{\text{psr}}$ are the Earth’s and the pulsar’s velocities, respectively. $v_{\text{screen}}$ is the velocity of the screen and $\mathbf{n}_{\text{screen}}$ is its 2D orientation vector in the sky.
@@
$v_{\text{sr}}=\left[s\mathbf{v}_{\text{Earth}}+(1-s)\mathbf{v}_{\text{psr}}\right]\cdot\mathbf{n}_{\text{screen}}-v_{\text{screen}}.$ (3)
In this case, $\mathbf{n}_{\text{screen}}$ stands for the main scattering direction and $v_{\text{screen}}$ is the screen velocity along this direction. We shall limit $\mathbf{n}_{\text{screen}}$ to be within [0, 180) degrees of a specified direction.
@@
$f_{\eta}\equiv\frac{\sqrt{2}v_{\text{sr}}}{\sqrt{s(1-s)d_{\text{psr}}}}.$ (5)
@@
Given the distance and the proper motion of the pulsar, we are left with three unknown parameters $s$, $v_{\text{screen}}$, and $\mathbf{n}_{\text{screen}}$. They are to be fitted from the annual modulation of the measurements $|f_{\eta}|(t)$.
II.1 Analytical understanding of the parameter dependence
II.1.1 1D scattering
Let us first consider the simpler case of 1D scattering, i.e., extremely anisotropic scattering. Approximating the Earth’s orbit as circular, $f_{\eta}(t)$ has a sinusoidal shape that is characterized by three quantities: the modulation amplitude $A_{\text{f}}\equiv(f_{\eta\text{, max}}-f_{\eta\text{, min}})/2$, the mean offset $\bar{f_{\eta}}$, and the yearly time $t_{\text{max}}$ when $f_{\eta}(t)$ reaches its maximum. They are dependent on the three parameters as follows.
When projected onto the direction of the pulsar, the Earth’s velocity vector $\mathbf{v}_{\text{Earth}}(t)$ forms an ellipse around the year (see Figure 1). This velocity ellipse is further projected onto the direction of $\mathbf{n}_{\text{screen}}$ to obtain $v_{\text{proj}}\equiv\mathbf{v}_{\text{Earth}}\cdot\mathbf{n}_{\text{screen}}$. The modulation amplitude $A_{\text{f}}$ is set by the maximum of $v_{\text{proj}}$,
$A_{\text{f}}=\sqrt{\frac{2s}{(1-s)}}\frac{v_{\text{proj,max}}}{\sqrt{d_{\text{psr}}}},$ (6)
and the maximum of $f_{\eta}(t)$ is reached at time $t_{\text{max}}$ when $|v_{\text{proj}}|=v_{\text{proj,max}}$. Therefore, from $t_{\text{max}}$ one can determine the orientation $\mathbf{n}_{\text{screen}}$. Given $\mathbf{n}_{\text{screen}}$, one can compute $v_{\text{proj,max}}$ and obtain s from the measurement of the modulation amplitude $A_{\text{f}}$.
@@ II.1 Analytical understanding of the parameter dependence
$\bar{f_{\eta}}=\left[1-\frac{1}{1-s}\frac{v_{\text{screen}}}{\mathbf{v}_{\text{psr}}\cdot\mathbf{n}_{\text{screen}}}\right]\sqrt{\frac{2(1-s)}{s}}\frac{\mathbf{v}_{\text{psr}}\cdot\mathbf{n}_{\text{screen}}}{{\sqrt{d_{\text{psr}}}}}.$ (7)
It depends on all three parameters. When the other two parameters are determined from $t_{\text{max}}$ and $A_{\text{f}}$, one can further determine $v_{\text{screen}}$ from $\bar{f_{\eta}}$.
Since a physical screen velocity is typically much smaller than the pulsar proper motion velocity, one generally has $v_{\text{screen}}/(\mathbf{v}_{\text{psr}}\cdot\mathbf{n}_{\text{screen}})/(1-s)\ll 1$, i.e.
$\bar{f_{n}}\approx\sqrt{\frac{2(1-s)}{s}}\frac{\mathbf{v}_{\text{psr}}\cdot\mathbf{n}_{\text{screen}}}{{\sqrt{d_{\text{psr}}}}}.$ (8)
@@
$A_{\text{f}}\bar{f_{\text{n}}}\approx\frac{2v_{\text{proj,max}}\mathbf{v}_{\text{psr}}\cdot\mathbf{n}_{\text{screen}}}{d_{\text{psr}}}\,$, (9)
@@
II.1.2 2D isotropic scattering
The full geometry of a general 2D scintillation screen cannot be constrained by the curvature measurements. It requires full modeling of the distribution of scattered power along and within a parabolic arc (Reardon et al., 2020), and is subject to degeneracies. The ideal case of 2D isotropic scattering, however, can be fully captured with a three-parameter model similar to the case of 1D scattering. In the early scintillation studies, scattering was thought to be produced by statistically isotropic ISM turbulence, and thus a 2D isotropic scattering was widely used. Deviation of the shape of $f_{\eta}(t)$ from a sinusoidal shape can be a sign of 2D scattering. For 2D isotropic scattering, the shape of $f_{\eta}(t)$ is in general not symmetric around the mean, but this asymmetric deviation from the sinusoidal shape is only strong when the velocity modulation by Earth is comparable to the yearly average of $v_{\text{sr}}/s$.
Apart from this strong-modulation case, the analytical framework for 1D scattering still holds approximately. The only difference is the projection of $\mathbf{v}_{\text{Earth}}$ to $v_{\text{proj}}$ that should be performed along the direction of $(1-s)\mathbf{v}_{\text{psr}}-v_{\text{screen}}\mathbf{n}_{\text{screen}}$ instead of $\mathbf{n}_{\text{screen}}$.
III Observation and Data Processing
III.1 Observation
@@
Table 1: The observed pulsars and their properties. $P$ (spin period) and $\dot{P}$ (spin-down rate) are retrieved from the ATNF catalogue (Manchester et al., 2005). All other properties of PSR B1237+25 and PSR B2021+51 were obtained from measurements using the NRAO Very Long Baseline Array (Brisken et al., 2002). The proper motion and the location of PSR B1842+14 were obtained from a program of observations carried out at Jodrell Bank to measure the proper motions of pulsars (Harrison et al., 1993). The DM of PSR B1842+14 was obtained from The Thousand-Pulsar-Array program on MeerKAT (Keith et al., 2024). Yao et al. (2020), using the YMW16 model (Yao et al., 2017), adopted an uncertainty of 40% of the distance (1.7 kpc) as a rough estimate for the one-sigma error of the DM-distance to PSR B1842+14. Therefore, we also adopt a distance of 1.7 ± 0.7 kpc for PSR B1842+14. The $\text{S}_{1400}$ of PSR B1842+14 was also obtained from the research by Yao et al. (2020).
Pulsar RAJ (hms) DECJ (dms) Parallax (mas) or Distance (kpc) PMRA (mas yr-1) PMDEC (mas yr-1) $P$ (ms) $\dot{P}$ (s s-1) $\text{S}_{1400}$ (mJy) DM ($\text{cm}^{-3}$ pc)
B1237+25 12:39:40.36 +24:53:50.01 $1.16\pm 0.08$ (mas) $106.82\pm 0.17$ $49.92\pm 0.18$ 1382 9.60 $\times 10^{-16}$ 20 9.28
B1842+14 18:44:54.88 +14:54:13.89 $1.7\pm 0.7$ (kpc) $9\pm 10$ $45\pm 6$ 375 $1.87\times 10^{-15}$ 1.8 41.49
B2021+51 20:22:49.87 +51:54:50.39 $0.50\pm 0.07$ (mas) $5.23\pm 0.17$ $11.54\pm 0.28$ 529 $3.06\times 10^{-15}$ 27 22.58
@@
III.2 Dynamic Spectra
We used DSPSR  [^1] https://psrchive.sourceforge.net/ (van Straten and Bailes, 2011) to fold the data in each frequency channel, with a sub-integration time chosen to be the closest value to 5 s that contains an integer number of pulse periods, and then summed the horizontal and vertical polarizations to obtain the total intensity. Next, we computed the average pulse profile in order to determine the on-pulse and off-pulse regions. The region below the median of the profile was identified as the off-pulse region, and the region above 10% of the peak amplitude with respect to the off-level was marked as the on-pulse region. For each sub-integration and frequency channel, we calculated the mean intensity in the on-pulse region minus the mean intensity in the off-pulse region. The result of this calculation is the dynamic spectrum.
We used the level of the off-pulse signal to identify the radio-frequency interference (RFI). After experimentation, we set the RFI threshold as: threshold = min + 1.5 × (median - min), where min and median are calculated from the off-pulse signals across all pixels. Any pixel exceeding this threshold is flagged as RFI.
Next, we applied Savitzky-Golay filters (Savitzky and Golay, 1964) to smooth the dynamic spectra. Finally, we applied Wiener filters (Wiener, 1949; Lin et al., 2021) to the dynamic spectra, filling the masked regions with the expected signal values obtained from the filter.